# Pi in the Sky: Neutron Stars with Exceptionally Light QCD Axions

### v2

@@ I Introduction
in the large $\mathcal{N}$ limit [44]. This motivates the study of lighter-than-expected QCD axions—hereafter referred to as exceptionally light QCD axions—that still solve the strong CP problem (for ${\mathcal{N}}$ odd). Even within this model, we note that as a small ${\theta}$ angle is itself technically natural, the very small $\varepsilon\ll 1$ parameter space, in particular $\varepsilon\lesssim 10^{-10}$, cannot be well-motivated. As we discuss below, due to their high density, neutron stars provide a unique environment to test exceptional axions close to the QCD line: values of $\varepsilon$ close to unity, and $\mathbb{Z}_{\mathcal{N}}$ models for modest values of $\mathcal{N}$.
@@
In this study, we focus on axion condensation in neutron stars for $\varepsilon\gtrsim 5\times 10^{-7}$. Earlier work has shown that for small axion masses ($m_{a}\lesssim 10^{-11}{\,\rm eV}$) the axion condensate extends far outside the radius of the star [36]. In this scenario, the extended axion field mediates a force between neutron stars and is constrained by LVK observations of neutron star mergers [36, 38, 39]. For a heavier axion with $m_{a}\gtrsim 10^{-8}{\,\rm eV}$, the axion field is largely confined inside the neutron star, and the effects of axion condensation on the neutron star interior need to be understood to predict the observational consequences. Prior work has considered neutron stars in the axion condensed phase in the non-interacting Fermi gas approximation [45].
@@
The outline of this paper is as follows. In Sec. II, we describe the most important features of axion condensation in a simple gas of non-interacting nucleons. In Sec. III, we discuss the effects of axion condensation including the potential effects of nuclear interactions, with a focus on ChiEFT and RMFT models. Section IV describes the impact that axion condensation has on the structure and properties of neutron stars within our nuclear model. In Sec. V, we describe novel nuclear density objects stabilized by an axion potential rather than gravity, first proposed in Refs. [37, 45] for a non-interacting Fermi gas. We present constraints in the $f_{a}$-$m_{a}$ parameter space for the phenomenological $\varepsilon$ model and the $Z_{\mathcal{N}}$ model based on neutron star observables in Sec. VI. Section VII concludes this manuscript with a summary and outlook.
@@ III Neutron and nuclear matter at $\theta=\pi$
C $0.16$ Matches the binding of neutron matter found using ChiEFT. See Figs. 2 & 4.
@@ III.2 Axion condensation in Chiral Effective Field Theory
Figure 2: Difference in the energy per particle (17) for the N2LO sim potentials [72] with two momentum cutoffs $\Lambda$ at $n_{B}=n_{\rm sat}$ as a function of $m_{\pi}$: $\Lambda=500$ MeV (left panel) and $\Lambda=450$ MeV (right panel), as annotated. In these potentials, we modify $m_{\pi}$, $f_{\pi}$, and the LO LEC $\tilde{C}_{{}^{1}S_{0}}$ (different line styles), whereas $g_{A}$ is kept at its physical value. The bands depict the variation of the LO LEC $\tilde{C}_{{}^{1}S_{0}}$ according to Eq. (19). The results were obtained at the Hartree-Fock level (green bands) and using second-order (blue bands) and third-order MBPT (red bands). See the main text for details.
@@
for two chiral potentials at $n_{B}=n_{\rm sat}$ as a function of $m_{\pi}$ (see annotations). Here, we use the family of bare N2LO potentials with chiral NN and 3N forces developed in Ref. [72]. These nonlocal potentials are called the sim potentials because they were simultaneously fit to scattering and bound-state observables in the pion-nucleon, NN, and few-nucleon sectors. Specifically, we use the two N2LO potentials with the NN and 3N momentum cutoffs $\Lambda=500$ (left panel) and $450$ MeV (right panel); both constrained by NN scattering data at laboratory energies $T_{\mathrm{lab}}\leqslant 290$ MeV [72]. To obtain the PNM energy per particle, we apply the Monte Carlo framework for MBPT nuclear matter calculations developed in Ref. [73]. The green bands in Fig. 2 depict the Hartree-Fock reference energy and the blue and red bands the MBPT results at second and third order, respectively. All MBPT calculations are performed at the normal-ordered two-body level. More details on MBPT for nuclear matter can be found in Ref. [48]. This MBPT framework uses the operatorial definition of the chiral potentials, allowing us to straightforwardly vary the values of physical constants, such as the pion mass, low-energy couplings, and more. Not all chiral potentials are given in their operatorial form; e.g., when evolved to lower resolution scales using renormalization group (RG) methods, their operatorial definition is usually unknown. For the solid lines in Fig. 2, we keep the values of the low-energy couplings at their best-fit values determined in Ref. [72] and vary only the pion mass, nucleon mass, and the pion decay constant. The $m_{\pi}$ dependence of the nucleon mass was discussed in Sec. II.1, and the $m_{\pi}$ dependence of the pion decay constant is given by
@@ III.3 Axion condensation in Relativistic Mean Field Theory
To derive concrete results for neutron stars, we use a model within RMFT (see Ref. [75] for a review). The Lagrangian of this theory is given by the following, omitting terms that will be discarded after taking the meson fields to have their mean field values.
$\begin{split}\mathcal{L}_{\rm RMF}&=\sum_{i=n,\,p}\overline{\psi}_{i}\left[i\not{\partial}-g_{\omega}\omega\gamma^{0}-\frac{1}{2}g_{\rho}\rho\gamma^{0}\tau_{3}-(m_{i}-g_{\sigma}\sigma)\right]\psi_{i}\\ &+\sum_{\ell=e,\,\mu}\overline{\psi}_{\ell}(i\not{\partial}-m_{\ell})\psi_{\ell}+\frac{1}{2}(-m_{\sigma}^{2}\sigma^{2}+m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2})+\mathcal{L}_{\sigma\omega\rho}\,.\end{split}$ (20)
$\mathcal{L}_{\sigma\omega\rho}$ contains all meson-meson couplings that are fit, along with the nucleon-meson couplings and meson masses, to known properties of nuclear matter. $\tau_{3}$ is the third Pauli matrix in isospin space. We use the IUFSU parameter set from Ref. [76]. The mean field values of the mesons and Fermi momenta of the nucleons and leptons are calculated by solving the Euler-Lagrange equations for each field and enforcing charge neutrality and $\beta$ equilibrium.
@@
For Scenario A, the simple case where only the nucleon masses are modified by finite $\theta$, Fig. 3 shows the energy per baryon, including axion energy and the reduction of the mass of the nucleons, of SNM, PNM, and beta equilibrated matter in this model. At fixed baryon density, the energy per baryon is minimized, leading to axion condensation due to the reduced nucleon mass. For small $\varepsilon$, the energy cost of having non-zero $\theta$ is small and axion condensation is favored at lower densities. Since the axions themselves contribute to the energy density, the energy per baryon of axion condensed matter will always diverge in the limit of zero density and a phase without axions will be favored. This has the effect of producing a minimum in the energy per baryon of axion condensed matter when $\varepsilon$ is sufficiently small and allows a stable zero pressure phase of nuclear matter at lower energy than normal nuclei. We will return to discuss objects made out of such matter in Sec. V.
@@
Not included thus far is the pion mass dependence of the nuclear force itself. The $\theta$-dependence of the masses of the $\sigma$ and $\rho$ mesons have been calculated in Refs. [56, 57]. They find that, comparing $\theta=0$ and $\theta=\pi$, the $\sigma$ mass decreases by about 6% and the $\rho$ mass decreases by about 2%. At low density, the RMFT EOS is dominated by the $\sigma$ meson and its modification is much larger than that of the vector mesons so we focus on the effect of modifying the scalar interaction. As a simple ansatz for modifications to the nuclear force, we modify the $\sigma$ mass with a tunable parameter $d_{\sigma}$ as follows:
@@
where $g_{\sigma}$ is a parameter of the RMFT model and $\chi$ is the volume fraction of nuclei when in a heterogeneous phase and unity in a homogeneous phase. At saturation density, $g_{\sigma}^{2}n_{\rm sat}/\sigma_{N}m_{\sigma}^{2}\simeq 11$ for the RMFT model we use so for Scenario B, axion condensation should be expected below $n_{\rm sat}$ for $\varepsilon\lesssim 0.8$ and in Scenario C, axion condensation should be expected below $n_{\rm sat}$, even for $\varepsilon=1$! The predictive power of this model is limited by the fact that we only consider isoscalar effects on the nuclear force. Axion condensation may also have an important impact on the symmetry energy of nuclear matter beyond the effect of $\Delta\sigma$. We expect this to have a more pronounced effect at high density where our model of the nucleon-axion interaction breaks down.
@@ III.4 EOS of axion condensed matter
We first focus on the left panel of Fig. 6. For $\varepsilon=10^{-6}$, the axion field itself contributes very little to the energy density and pressure at the densities pictured, and Scenario A looks very similar to the case with no axions. At pressures of order $\varepsilon m_{\pi}^{2}f_{\pi}^{2}\simeq 10\times\varepsilon\,{\rm MeV/fm}^{3}$ the curve for Scenario A will diverge from the EOS without axions and go off the bottom of the plot as in the center panel. This occurs because zero pressure in the axion condensed phase occurs when the pressure of nuclear matter exactly cancels the negative axion pressure, corresponding to a much larger baryon density than normally found in nuclear matter at zero pressure. The point where the curves for Scenario B and C in the left panel diverge from Scenario A and the EOS without axions corresponds to the neutron drip phase transition for Scenario A and the EOS without axions. When $K_{\sigma}$ is large enough that the inner crust is absent, the pressure is larger (the EOS is “stiffer”) at energy densities where ordinarily dripped neutrons would be present because there are fewer degrees of freedom in the system, resulting in the lepton density being larger. At the crust-core transition, however, there is a large jump in energy density from the mixed phase to bulk nuclear matter, resulting in a lower pressure (the EOS is “softer”) in the outermost part of the core.
@@ IV.1 Neutron star structure in axion condensed stars
Although our treatment of axion condensed matter does not apply at arbitrarily high density, we can still speculate about the effects of axion condensation on neutron star masses and radii. A similar calculation was performed in Ref. [45] for a non-interacting gas of neutrons. As can be seen in Fig. 6, at low density, the EOS is typically softer for $\theta=\pi$ and $\varepsilon$ not too small because of the contribution of the axion field. A naive extrapolation of nuclear matter to high density with only the axion-nucleon interaction we have considered thus far will eventually yield a stiffer EOS because of the lighter nucleon mass. This results in heavy neutron stars having larger radii and larger maximum mass. Figure 9 shows the relationship between mass and radius for neutron stars with various $\varepsilon$ within this simplistic picture of the high density EOS and treating the domain wall as thin ($1/m_{a}\ll r_{\rm NS})$.
@@
For $\varepsilon$ where the lowest density parts of the crust are missing, low mass stars may have very small radii, with a mass-radius relation akin to strange quark stars. Unlike a strange quark star, however, neutron stars with a condensed axion field lack a confining potential at their surface, and as a result have a thin crust of matter at intermediate $\theta$ in the domain wall region. Since the density and chemical potential smoothly go to zero on a length scale determined by the axion field profile, unusual surface properties of strange quark stars (see, e.g., Ref. [79] for a detailed discussion) are not present in axion condensed stars.

### v3

# Neutron Stars with Exceptionally Light QCD Axions

@@ I Introduction
in the large $\mathcal{N}$ limit [44]. This motivates the study of lighter-than-expected QCD axions—hereafter referred to as exceptionally light QCD axions—that still solve the strong CP problem (for ${\mathcal{N}}$ odd). Even within this model, we note that as a small ${\theta}$ angle is itself technically natural, the very small $\varepsilon\ll 1$ parameter space, in particular $\varepsilon\lesssim 10^{-10}$, cannot be well-motivated: it would introduce a tuning at the level of the original tuning of the Strong CP problem. As we discuss below, due to their high density, neutron stars provide a unique environment to test exceptional axions close to the QCD line: values of $\varepsilon$ close to unity, and $\mathbb{Z}_{\mathcal{N}}$ models for modest values of $\mathcal{N}$. We note that the vacuum structure of the $\mathbb{Z}_{\mathcal{N}}$ model is significantly different from the phenomenological $\varepsilon$ model and can result in different dynamics and constraints.
@@
In this study, we focus on axion condensation in neutron stars for $\varepsilon\gtrsim 5\times 10^{-7}$. Earlier work has shown that for small axion masses ($m_{a}\lesssim 10^{-11}{\,\rm eV}$) the axion condensate extends far outside the radius of the star [36]. In this scenario, the extended axion field mediates a force between neutron stars and is constrained by LVK observations of neutron star mergers [36, 38, 39]. For a heavier axion with $m_{a}\gtrsim 10^{-8}{\,\rm eV}$, the axion field is largely confined inside the neutron star, and the effects of axion condensation on the neutron star interior need to be understood to predict the observational consequences. Prior work has considered neutron stars in the axion condensed phase in the non-interacting Fermi gas approximation [45] and in hybrid stars where axion interactions were included in the quark phase (but not the nuclear phase) via an effective ’t Hooft determinant interaction in the NJL model [46].
@@
The outline of this paper is as follows. In Sec. II, we describe the most important features of axion condensation in a simple gas of non-interacting nucleons. In Sec. III, we discuss the effects of axion condensation including the potential effects of nuclear interactions, with a focus on ChiEFT and RMFT models. Section IV describes the impact that axion condensation has on the structure and properties of neutron stars within our nuclear model. In Sec. V, we describe novel nuclear density objects stabilized by an axion potential rather than gravity, first proposed in Refs. [37, 45] for a non-interacting Fermi gas. We present constraints in the $f_{a}$-$m_{a}$ parameter space for the phenomenological $\varepsilon$ model and comment on the $\mathbb{Z}_{\mathcal{N}}$ model based on neutron star observables in Sec. VI. Section VII concludes this manuscript with a summary and outlook.
@@ III Neutron and nuclear matter at $\theta=\pi$
C $0.16$ Matches the largest binding of neutron matter found using ChiEFT. See Figs. 2 & 4.
@@ III.2 Axion condensation in Chiral Effective Field Theory
Figure 2: Difference in the energy per particle (17) for the N2LO sim potentials [73] with two momentum cutoffs $\Lambda$ at $n_{B}=n_{\rm sat}$ as a function of $m_{\pi}$: $\Lambda=500$ MeV (left panel) and $\Lambda=450$ MeV (right panel), as annotated. In these potentials, we modify $m_{\pi}$, $f_{\pi}$, and the LO LEC $\tilde{C}_{{}^{1}S_{0}}$ (different line styles), whereas $g_{A}$ is kept at its physical value. The bands depict the variation of the LO LEC $\tilde{C}_{{}^{1}S_{0}}$ according to Eq. (19). The results were obtained at the Hartree-Fock level (green bands) and using second-order (blue bands) and third-order MBPT (red bands). The energy per particle at the physical pion mass obtained from on our third-order calculations is $E(n_{0})=16.94\,\mathrm{MeV}$ and $E(n_{0})=17.21\,\mathrm{MeV}$ for $\Lambda=450\,\mathrm{MeV}$ and $\Lambda=500\,\mathrm{MeV}$, respectively. See the main text for details.
@@
for two chiral potentials at $n_{B}=n_{\rm sat}$ as a function of $m_{\pi}$ (see annotations). Here, we use the family of bare N2LO potentials with chiral NN and 3N forces developed in Ref. [73]. These nonlocal potentials are called the sim potentials because they were simultaneously fit to scattering and bound-state observables in the pion-nucleon, NN, and few-nucleon sectors. Specifically, we use the two N2LO potentials with the NN and 3N momentum cutoffs $\Lambda=500$ (left panel) and $450$ MeV (right panel); both constrained by NN scattering data at laboratory energies $T_{\mathrm{lab}}\leqslant 290$ MeV [73]. The analytic expressions that define these potentials and were used in this work are summarized in the Appendix of Ref. [74] and Section 5.1.2 in Ref. [75]. To obtain the PNM energy per particle, we apply the Monte Carlo framework for MBPT nuclear matter calculations developed in Ref. [76]. The green bands in Fig. 2 depict the Hartree-Fock reference energy and the blue and red bands the MBPT results at second and third order, respectively. All MBPT calculations are performed at the normal-ordered two-body level. More details on MBPT for nuclear matter can be found in Ref. [49]. This MBPT framework uses the operatorial definition of the chiral potentials, allowing us to straightforwardly vary the values of physical constants, such as the pion mass, low-energy couplings, and more. Not all chiral potentials are given in their operatorial form; e.g., when evolved to lower resolution scales using renormalization group (RG) methods, their operatorial definition is usually unknown. For the solid lines in Fig. 2, we keep the values of the low-energy couplings at their best-fit values determined in Ref. [73] and vary only the pion mass, nucleon mass, and the pion decay constant. The $m_{\pi}$ dependence of the nucleon mass was discussed in Sec. II.1, and the $m_{\pi}$ dependence of the pion decay constant is given by
@@ III.3 Axion condensation in Relativistic Mean Field Theory
To derive concrete results for neutron stars, we use a model within RMFT (see Ref. [78] for a review). The Lagrangian of this theory is given by the following, omitting terms that will be discarded after taking the meson fields to have their mean field values,
$\begin{split}\mathcal{L}_{\rm RMF}&=\sum_{i=n,\,p}\overline{\psi}_{i}\left[i\not{\partial}-g_{\omega}\omega\gamma^{0}-\frac{1}{2}g_{\rho}\rho\gamma^{0}\tau_{3}-(m_{i}-g_{\sigma}\sigma)\right]\psi_{i}\\ &+\sum_{\ell=e,\,\mu}\overline{\psi}_{\ell}(i\not{\partial}-m_{\ell})\psi_{\ell}+\frac{1}{2}(-m_{\sigma}^{2}\sigma^{2}+m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2})+\mathcal{L}_{\sigma\omega\rho}\,,\end{split}$ (20)
where the nucleon-meson couplings are denoted by $g_{\sigma},g_{\omega}$ and $g_{\rho}$, the meson masses by $m_{\sigma},m_{\omega}$ and $m_{\rho}$, and $m_{i}$ are the nucleon masses. The term $\mathcal{L}_{\sigma\omega\rho}$ contains all meson-meson couplings that are fit, along with the nucleon-meson couplings to known properties of nuclear matter for the canonical choice of meson masses $m_{\sigma}=550$ MeV and $m_{\omega}=m_{\rho}=770$ MeV. $\tau_{3}$ is the third Pauli matrix in isospin space. We use the IUFSU parameter set from Ref. [79] because it is optimized for the calculation of finite nuclei and our study will focus on the crust. The mean field values of the mesons and Fermi momenta of the nucleons and leptons are calculated by solving the Euler-Lagrange equations for each field and enforcing charge neutrality and $\beta$ equilibrium.
Since mean field models do not explicitly include pion fields, incorporating the effects of a reduced pion mass on nuclear interactions presents a challenge. To address this, we model the increased attraction predicted by ChiEFT at lower pion masses by reducing the mass of the $\sigma$ meson. This approach is motivated by the fact that the two-pion exchange potential—which provides significant attraction in ChiEFT—can be effectively represented by scalar ($\sigma$) meson exchange in mean field models.
We consider three scenarios in our analysis. In Scenario A, we assume no change in the nuclear interaction, meaning the $\sigma$-meson mass is independent of $\theta$. In Scenarios B and C, we introduce a modest reduction in the $\sigma$-meson mass to mimic the decrease in energy per particle predicted by ChiEFT at $m_{\pi}=82$ MeV. The specific values of $m_{\sigma}$ used in these scenarios are informed by studies of the $\theta$ dependence of meson masses reported in Refs.[57, 58].
@@
As noted earlier, Scenario A is the simple case where interactions are unmodified and only the nucleon masses are modified by finite $\theta$. For this case, Fig. 3 shows the energy per baryon, including axion energy and the reduction of the mass of the nucleons, of SNM, PNM, and beta equilibrated matter in this model. At fixed baryon density, the energy per baryon is minimized, leading to axion condensation due to the reduced nucleon mass. For small $\varepsilon$, the energy cost of having non-zero $\theta$ is small and axion condensation is favored at lower densities. Since the axions themselves contribute to the energy density, the energy per baryon of axion condensed matter will always diverge in the limit of zero density and a phase without axions will be favored. This has the effect of producing a minimum in the energy per baryon of axion condensed matter when $\varepsilon$ is sufficiently small and allows a stable zero pressure phase of nuclear matter at lower energy than normal nuclei. We will return to discuss objects made out of such matter in Sec. V.
@@
Not included thus far is the pion mass dependence of the nuclear force itself. As noted earlier, we use a simple ansatz for modifications to the nuclear force, we modify the $\sigma$ mass with a tunable parameter $d_{\sigma}$ as follows:
@@
where $g_{\sigma}$ is a parameter of the RMFT model and $\chi$ is the volume fraction of nuclei when in a heterogeneous phase and unity in a homogeneous phase. At saturation density, $g_{\sigma}^{2}n_{\rm sat}/\sigma_{N}m_{\sigma}^{2}\simeq 11$ for the RMFT model we use so for Scenario B, axion condensation should be expected below $n_{\rm sat}$ for $\varepsilon\lesssim 0.8$ and in Scenario C, axion condensation should be expected below $n_{\rm sat}$, even for $\varepsilon=1$! While this is a striking prediction, it should be noted that RMFT is a model for the nuclear interaction that does not have explicit pions and we are parameterizing the expected effects of a reduced pion mass via the meson masses and where Scenario C is chosen to recreate the most extreme prediction of ChiEFT. While the $\sigma$ meson is typically understood as a two pion resonance, this comparison is certainly incomplete. More detailed study is required to determine the fate of neutron matter at saturation density. The predictive power of this model is additionally limited by the fact that we only consider isoscalar effects on the nuclear force. Axion condensation may also have an important impact on the symmetry energy of nuclear matter beyond the effect of $\Delta\sigma$. We expect this to have a more pronounced effect at high density where our model of the nucleon-axion interaction breaks down.
@@ III.4 EOS of axion condensed matter
We first focus on the left panel of Fig. 6. For $\varepsilon=10^{-6}$, the axion field itself contributes very little to the energy density and pressure at the densities pictured, and Scenario A looks very similar to the case with no axions. At pressures of order $\varepsilon m_{\pi}^{2}f_{\pi}^{2}\simeq 10\times\varepsilon\,{\rm MeV/fm}^{3}$ the curve for Scenario A will diverge from the EOS without axions and go to zero pressure at finite energy density. For $\varepsilon=10^{-6}$ this corresponds to $p\simeq 10^{-5}\rm{MeV/fm}^{3}$, outside the plot range shown. This occurs because zero pressure in the axion condensed phase occurs when the pressure of nuclear matter exactly cancels the negative axion pressure, corresponding to a much larger baryon density than normally found in nuclear matter at zero pressure. The point where the curves for Scenario B and C in the left panel diverge from Scenario A and the EOS without axions corresponds to the neutron drip phase transition for Scenario A and the EOS without axions. When $K_{\sigma}$ is large enough that the inner crust is absent, the pressure is larger (the EOS is “stiffer”) at energy densities where ordinarily dripped neutrons would be present because there are fewer degrees of freedom in the system, resulting in the lepton density being larger. At the crust-core transition, however, there is a large jump in energy density from the mixed phase to bulk nuclear matter, resulting in a lower pressure (the EOS is “softer”) in the outermost part of the core.
@@ IV.1 Neutron star structure in axion condensed stars
Although our treatment of axion condensed matter does not apply at arbitrarily high density, we can still speculate about the effects of axion condensation on neutron star masses and radii. A similar calculation was performed in Ref. [45] for a non-interacting gas of neutrons, see also Ref. [82] for scalar effects on the crust thickness and EoS. As can be seen in Fig. 6, at low density, the EOS is typically softer for $\theta=\pi$ and $\varepsilon$ not too small because of the contribution of the axion field. A naive extrapolation of nuclear matter to high density with only the axion-nucleon interaction we have considered thus far will eventually yield a stiffer EOS because of the lighter nucleon mass. This results in heavy neutron stars having larger radii and larger maximum mass. Figure 9 shows the relationship between mass and radius for neutron stars with various $\varepsilon$ within this simplistic picture of the high density EOS and treating the domain wall as thin ($1/m_{a}\ll r_{\rm NS})$.
@@
For $\varepsilon$ where the lowest density parts of the crust are missing, low mass stars may have very small radii, with a mass-radius relation akin to strange quark stars. Unlike a strange quark star, however, neutron stars with a condensed axion field lack a confining potential at their surface, and as a result have a thin crust of matter at intermediate $\theta$ in the domain wall region. Since the density and chemical potential smoothly go to zero on a length scale determined by the axion field profile, unusual surface properties of strange quark stars (see, e.g., Ref. [83] for a detailed discussion) are not present in axion condensed stars. For the same reason, the radius of an axion condensed star does not go to zero for vanishing mass for any $\varepsilon$ but will always be limited by the finite size of the domain wall. The minimum stable mass and radius of a compact axion condensed object will be addressed in greater detail in Sec. V.